MINUTES FROM THE BOOSTER SUPERVISOR MEETING
DATE: August 20, 2002
PSB home page, PSB
PRESENT: M. Benedikt, C. Carli,
G. Cyvoct, K. Hanke, C. Hills, D. KŸchler, M. O'Neil
Supervisor Linac: K. Hanke ˆ
Supervisor PSB: M. Chanel ˆ C. Carli
- Problems with the
electronics of the beam transformer TRA6 have been solved by F. Lenardon
- Beam losses have been
observed. They may have been caused by quadrupoles (LA2.QDN4S and
LA2.QDN10S) not working correctly (power supply did not follow AQN values).
The problem with the quadrupole power supplies has been resolved by the
exchange of a AQN card. However it is not clear, whether if similar losses
have been observed after the intervantion.
- A (single) source flashover
caused the vacuum valve LI.VVS10 to shut.
- Fluctuations of the magnetic
field at the flat-top (up to 10 G at 1.0 GeV and 6 G at
1.4 GeV) are still present. A first intervention (exchange of
regulation electronics of the MPS) did not resolve. One hypothesis to
explain this jitter is a problem at an integration to convert the time
derivative of the current (which is programmed and provided during the
cycle by a GFA) to provide the current, needed as input to the MPS
electronics. This would imply that those fluctuations are present since
along time (at least one year) and have only been observed now.
- Baseline drift (with high
intensity) of the pick-ups BT.UES30 and BT.UES40 in the PSB transfer lines.
In order to solve this old problem, the circuits at the entrance to the amplifiers
have been modified (capacity plus some resistances added to filter low
frequencies out) by M. Gasior. The behaviour has been improved.
- BSM (Bunch Shape Measurement)
does not work properly (e.g. sometimes measurements were executed at low
energy, while asking for a measurement at high energy via the application
program). This is caused by the state of the scope doing the measurement.
For the future, to be able to solve such problems quickly from the MCR,
additional parameters (e.g. choice of trigger- internal/external and
trigger level) should be added to the “Specialist page” of the BSM
- Ions : PSB will start the
set-up with ion beams on September 3, 2002. Thus, tests without beam
(e.g. RF, distributors …) should take place during the week before, i.e.
starting on August 27.
- Next BS meeting Tuesday September 3, 2002
at 10.00 in the MCR Glassbox.
Minutes by Christian Carli.